OpenStudy (anonymous):

Stacy and Travis are rock climbing. Stacy's rope is 4 feet shorter than 3 times the length of Travis' rope. Stacy's rope is 23 feet long. Write and solve an equation to find the length t of Travis' rope.

OpenStudy (texaschic101):

s = 23 s = 3t - 4 s + 4 = 3t (s + 4) / 3 = t (23 + 4) / 3 = t 27/3 = t 9 = t check.. s = 3t - 4 23 = 3(9) - 4 23 = 27 - 4 23 = 23 (correct) travis's rope is 9 ft